# Simplifying Debiased Inference via Automatic Differentiation and Probabilistic Programming

### v1

@@ D.4.6   Bounded affine map
D.4.7   Constant map
Fix $c$ in a Hilbert space $\mathcal{W}$. Let $\zeta$ be the constant map that takes as input an element $u$ belonging to a subset $\mathcal{U}$ of a Hilbert space $\mathcal{T}$ and, regardless of form $u$ takes, returns $c$. This map is a special case of the affine maps considered in Appendix D.4.6 with $\kappa$ equal to a zero operator. Hence, $\zeta$ is Hadamard differentiable with $\dot{\zeta}_{u}(t)=0$ and $\dot{\zeta}_{u}^{*}(w)=0$.
D.4.8   Coordinate projection
Let $\mathcal{U}=\mathcal{R}_{1}\oplus\mathcal{R}_{2}$, where $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are Hilbert spaces. Define $\zeta$ so that $\zeta(u)=u_{1}$, where $u=(u_{1},u_{2})$. Note that $\zeta$ is a linear map. It is also bounded since, for all $u=(u_{1},u_{2})\in\mathcal{U}$, $\|\zeta(u)\|_{\mathcal{R}_{1}}=\|u_{1}\|_{\mathcal{R}_{1}}\leq\|u\|_{\mathcal{R}_{1}\oplus\mathcal{R}_{2}}$. Hence, by the results in Appendix D.4.6, $\zeta$ is Hadamard differentiable with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$. To see that $\zeta^{*}(w)=(w,0)$, observe that, for any $w\in\mathcal{R}_{1}$ and $u\in\mathcal{U}$, $\langle\zeta(u),w\rangle_{\mathcal{R}_{1}}=\langle u_{1},w\rangle_{\mathcal{R}_{1}}=\langle u,(w,0)\rangle_{\mathcal{R}_{1}\oplus\mathcal{R}_{2}}$.
D.4.9   Change of measure
@@
We establish the Hadamard differentiability of $\zeta$ by showing that it is a special case of the affine maps considered in Appendix D.4.6. Clearly $\zeta$ is linear. To see that it is bounded, note that, for all $[u]_{1}\in L^{2}(\lambda_{1})$,
@@
where the essential supremum norm on the right-hand side is finite by assumption. Hence, by Appendix D.4.6, $\zeta$ is Hadamard differentiable at any $u\in\mathcal{U}$ with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$.
@@ D.4.10   Lifting to new domain
We establish the Hadamard differentiability of $\zeta$ by showing that it is a special case of the affine maps considered in Appendix D.4.6. Clearly $\zeta$ is linear. It is also bounded since, for any $u\in L^{2}(Q_{X})$, $\|\zeta(u)\|_{L^{2}(Q)}=\|u\|_{L^{2}(Q_{X})}$. Hence, by Appendix D.4.6, $\zeta$ is Hadamard differentiable at any $u\in\mathcal{U}$ with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$.
@@ D.4.11   Fix binary argument
To show $\zeta$ is Hadamard differentiable at a generic $u\in L^{2}(Q_{A,X})$, we express it as a composition $\zeta_{2}\circ\zeta_{1}$ of Hadamard differentiable maps and then apply the chain rule. The first of these maps is the change of measure $\zeta_{1}:L^{2}(\lambda_{1})\rightarrow L^{2}(\lambda_{2})$ studied in Appendix D.4.9, where $\lambda_{1}=Q_{A,X}$ and $\lambda_{2}=Q_{A=1,X}$ with $Q_{A=1,X}$ the distribution on $\{0,1\}\times\mathcal{X}$ with Radon-Nikodym derivative $\frac{dQ_{A=1,X}}{dQ_{A,X}}(a,x)=\frac{a}{\pi(x)}$. This map is Hadamard differentiable with $\dot{\zeta}_{1,u}(t)(a,x)=t(a,x)$ and $\dot{\zeta}_{1,u}^{*}(w)(a,x)=\frac{a}{\pi(x)}w(a,x)$. The second is the map $\zeta_{2}:L^{2}(Q_{A=1,X})\rightarrow L^{2}(Q_{X})$ defined so that $\zeta_{2}(\tilde{u})(x)=\tilde{u}(1,x)$. This is a bounded linear map, so by Appendix D.4.6 it is Hadamard differentiable with $\dot{\zeta}_{2,u}=\zeta_{2}$ and $\dot{\zeta}_{2,u}^{*}=\zeta_{2}^{*}$. It is also straightforward to verify that $\zeta_{2}^{*}(w)(a,x)=w(x)$. Since $\dot{\zeta}_{u}^{*}$ is the adjoint of a composition, we also have that $\dot{\zeta}_{u}^{*}(w)=\dot{\zeta}_{1,u}^{*}\circ\dot{\zeta}_{2,\zeta_{1}(u)}^{*}(w)$. Plugging in the values for the differential operators and adjoints of $\zeta_{1},\zeta_{2}$, this yields that $\dot{\zeta}_{u}(t)(x)=t(1,x)$ and $\dot{\zeta}_{u}^{*}(w)(a,x)=\frac{a}{\pi(x)}w(x)$.
D.5   Primitives for marginal quantities and semiparametric models
@@
The primitives in Table 1 can be readily adapted to semiparametric models. This can be done by using that, for any semiparametric $\mathcal{M}^{\prime}\subset\mathcal{M}$, the restriction of a totally pathwise differentiable primitive $\theta:\mathcal{M}\times\mathcal{U}\rightarrow\mathcal{W}$ to ×$\mathcal{M}^{\prime}\times\mathcal{U}$ is itself totally pathwise differentiable. The adjoint of this new primitive at $(P,u)$ is the same as that of $\theta$, except its first entry is replaced by its orthogonal projection onto the tangent space ˙$\dot{\mathcal{M}}_{P}^{\prime}$. Thus, when $\psi:\mathcal{M}^{\prime}\rightarrow\mathbb{R}$ can be expressed as a composition of restrictions of primitives in Table 1, Algorithm 2 returns its semiparametric EIF relative to $\mathcal{M}^{\prime}$. Naturally, the returned form of the EIF will be easiest to work with when the projection onto ˙$\dot{\mathcal{M}}_{P}^{\prime}$ has a known closed form.
Appendix E Step-by-step illustrations of Algorithm 2
Like Table 2 from the main text, Tables S2 and S3 provide step-by-step illustrations of Algorithm 2. The first studies the expected density parameter, and the second studies an expected conditional covariance parameter. The primitives used to express these parameters are detailed in the table captions. The value of $f_{0}$ in the bottom row of each table is the EIF at a distribution $P$ in a nonparametric model $\mathcal{M}$. In these examples, $Q\in\mathcal{M}$ is a fixed distribution used to define the ambient Hilbert spaces of some of the primitives used to represent $\psi$. We suppose $Q$ is mutually absolutely continuous with $P$ and let $\frac{p}{q}:=\frac{dP}{dQ}$ and $\frac{p_{X}}{q_{X}}:=\frac{dP_{X}}{dQ_{X}}$.
Table S2: Step-by-step evaluation of Algorithm 2’s computations of the EIF of the expected density $\psi(\tilde{P})=\int\tilde{p}(z)^{2}\lambda(dz)$, where $\tilde{p}=d\tilde{P}/d\lambda$ for a $\sigma$-finite measure $\lambda$. Under regularity conditions, the parameter $\psi$ is expressible as a composition of a root-density $\theta_{1}(\tilde{P},0)=\tilde{p}^{1/2}\in L^{2}(\lambda)$, pointwise square $\theta_{2}(\tilde{P},h_{1})=h_{1}^{2}\in L^{2}(\lambda)$, change of measure $\theta_{3}(\tilde{P},h_{2})=h_{2}\in L^{2}(Q)$, and mean operator $\theta_{4}(\tilde{P},h_{3})=E_{\tilde{P}}[h_{3}(Z)]\in\mathbb{R}$. The value of $f_{0}$ in the bottom row is the efficient influence function of $\psi$ at $P$.
@@ 0 $\bm{2[p-\psi(P)]}$
This lemma is a direct consequence of Lemma S5 and Appendix D.3.1 and so the proof is omitted.
Table S3: Step-by-step evaluation of Algorithm 2’s computations of the EIF of the expected conditional covariance $\psi(\tilde{P})=E_{\tilde{P}}[\mathrm{Cov}_{\tilde{P}}(A,Y\,|\,X)]$, where $Z=(X,A,Y)$. Let $\mathscr{C}:z\mapsto x$. Under regularity conditions, the parameter $\psi$ is expressible as a composition of conditional means $\theta_{1}(\tilde{P},0)(\cdot)=E_{\tilde{P}}[Y\,|\,X=\cdot\,]\in L^{2}(Q_{X})$ and $\theta_{2}(\tilde{P},0)(\cdot)=E_{\tilde{P}}[A\,|\,X=\cdot\,]\in L^{2}(Q_{X})$, liftings $\theta_{3}(\tilde{P},h_{1})=(z\mapsto h_{1}(x))\in L^{2}(Q)$ and $\theta_{4}(\tilde{P},h_{2})=(z\mapsto h_{2}(x))\in L^{2}(Q)$, a pointwise operation $\theta_{5}(\tilde{P},(h_{3},h_{4}))=(z\mapsto ay-h_{3}(z)h_{4}(z))\in L^{2}(Q)$, and a mean operator $\theta_{6}(\tilde{P},h_{5})=E_{\tilde{P}}[h_{5}(Z)]\in\mathbb{R}$. Below $\mu_{P}^{Y}(x):=E_{P}[Y\,|\,X=x]$ and $\mu_{P}^{A}(x):=E_{P}[A\,|\,X=x]$. The value of $f_{0}$ in the bottom row is the efficient influence function of $\psi$ at $P$. Other expressions of this parameter, such as ones using the conditional covariance primitive studied in Appendix D.2.3, can also be used to compute the EIF.
@@ 0 $\bm{z\mapsto[a-\mu_{P}^{A}(x)][y-\mu_{P}^{Y}(x)]-\psi(P)}$
Appendix F Study of remainder in von Mises expansion of nonparametric $R^{2}$
@@ 7 $-\frac{1}{\widehat{h}_{2}}$ 00 00 00 -- $\frac{\widehat{h}_{7}}{\widehat{h}_{2}^{2}}$- 00
Forward and backward pass nuisances are estimated as described in Appendix G. In the forward routines, this entails setting $\widehat{h}_{1}:z\mapsto y$, $\widehat{h}_{2}=\mathrm{Var}_{P_{\mathscr{I}}}(Y)$ with $P_{\mathscr{I}}$ the empirical distribution of $Z_{\mathscr{I}}:=\{Z_{i}:i\in\mathscr{I}\}$, $\widehat{h}_{3}:z\mapsto y$, $\widehat{h}_{4}:x\mapsto\widehat{E}_{4}[Y\,|\,X=x]$ with the conditional mean estimate returned by a nonparametric regression procedure, $\widehat{h}_{5}:z\mapsto\widehat{h}_{4}(x)$, $\widehat{h}_{6}:z\mapsto[y-\widehat{h}_{5}(z)]^{2}$, $\widehat{h}_{7}=E_{P_{\mathscr{I}}}[\widehat{h}_{6}(Z)]$, and $\widehat{h}_{8}=1-\widehat{h}_{7}/\widehat{h}_{2}$. In the backward routines, the only nuisances that require estimation are $E_{P}[Y\,|\,X=\,\cdot\,]$ and $E_{P}[Y]$, which arise when obtaining the estimates $\widehat{\vartheta}_{5}$ and $\widehat{\vartheta}_{2}$, respectively. We denote these estimates by $\widehat{E}_{5}[Y\,|\,X=\,\cdot\,]$ and $\widehat{E}_{2}[Y]$, respectively.
@@
Appendix G Nuisance estimation routines for primitives in Table 1
G.1   Overview
@@
G.2   Maps depend nontrivially on both of their arguments
G.2.1   Conditional mean

### v3

@@ C.4.6   Bounded affine map
C.4.7   Constant map
Fix $c$ in a Hilbert space $\mathcal{W}$. Let $\zeta$ be the constant map that takes as input an element $u$ belonging to a subset $\mathcal{U}$ of a Hilbert space $\mathcal{T}$ and, regardless of form $u$ takes, returns $c$. This map is a special case of the affine maps considered in Appendix C.4.6 with $\kappa$ equal to a zero operator. Hence, $\zeta$ is Hadamard differentiable with $\dot{\zeta}_{u}(t)=0$ and $\dot{\zeta}_{u}^{*}(w)=0$.
C.4.8   Coordinate projection
Let $\mathcal{U}=\mathcal{R}_{1}\oplus\mathcal{R}_{2}$, where $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are Hilbert spaces. Define $\zeta$ so that $\zeta(u)=u_{1}$, where $u=(u_{1},u_{2})$. Note that $\zeta$ is a linear map. It is also bounded since, for all $u=(u_{1},u_{2})\in\mathcal{U}$, $\|\zeta(u)\|_{\mathcal{R}_{1}}=\|u_{1}\|_{\mathcal{R}_{1}}\leq\|u\|_{\mathcal{R}_{1}\oplus\mathcal{R}_{2}}$. Hence, by the results in Appendix C.4.6, $\zeta$ is Hadamard differentiable with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$. To see that $\zeta^{*}(w)=(w,0)$, observe that, for any $w\in\mathcal{R}_{1}$ and $u\in\mathcal{U}$, $\langle\zeta(u),w\rangle_{\mathcal{R}_{1}}=\langle u_{1},w\rangle_{\mathcal{R}_{1}}=\langle u,(w,0)\rangle_{\mathcal{R}_{1}\oplus\mathcal{R}_{2}}$.
C.4.9   Change of measure
@@
We establish the Hadamard differentiability of $\zeta$ by showing that it is a special case of the affine maps considered in Appendix C.4.6. Clearly $\zeta$ is linear. To see that it is bounded, note that, for all $[u]_{1}\in L^{2}(\lambda_{1})$,
@@
where the essential supremum norm on the right-hand side is finite by assumption. Hence, by Appendix C.4.6, $\zeta$ is Hadamard differentiable at any $u\in\mathcal{U}$ with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$.
@@ C.4.10   Lifting to new domain
We establish the Hadamard differentiability of $\zeta$ by showing that it is a special case of the affine maps considered in Appendix C.4.6. Clearly $\zeta$ is linear. It is also bounded since, for any $u\in L^{2}(Q_{X})$, $\|\zeta(u)\|_{L^{2}(Q)}=\|u\|_{L^{2}(Q_{X})}$. Hence, by Appendix C.4.6, $\zeta$ is Hadamard differentiable at any $u\in\mathcal{U}$ with $\dot{\zeta}_{u}=\zeta$ and $\dot{\zeta}_{u}^{*}=\zeta^{*}$.
@@ C.4.11   Fix binary argument
To show $\zeta$ is Hadamard differentiable at a generic $u\in L^{2}(Q_{A,X})$, we express it as a composition $\zeta_{2}\circ\zeta_{1}$ of Hadamard differentiable maps and then apply the chain rule. The first of these maps is the change of measure $\zeta_{1}:L^{2}(\lambda_{1})\rightarrow L^{2}(\lambda_{2})$ studied in Appendix C.4.9, where $\lambda_{1}=Q_{A,X}$ and $\lambda_{2}=Q_{A=1,X}$ with $Q_{A=1,X}$ the distribution on $\{0,1\}\times\mathcal{X}$ with Radon-Nikodym derivative $\frac{dQ_{A=1,X}}{dQ_{A,X}}(a,x)=\frac{a}{\pi(x)}$. This map is Hadamard differentiable with $\dot{\zeta}_{1,u}(t)(a,x)=t(a,x)$ and $\dot{\zeta}_{1,u}^{*}(w)(a,x)=\frac{a}{\pi(x)}w(a,x)$. The second is the map $\zeta_{2}:L^{2}(Q_{A=1,X})\rightarrow L^{2}(Q_{X})$ defined so that $\zeta_{2}(\tilde{u})(x)=\tilde{u}(1,x)$. This is a bounded linear map, so by Appendix C.4.6 it is Hadamard differentiable with $\dot{\zeta}_{2,u}=\zeta_{2}$ and $\dot{\zeta}_{2,u}^{*}=\zeta_{2}^{*}$. It is also straightforward to verify that $\zeta_{2}^{*}(w)(a,x)=w(x)$. Since $\dot{\zeta}_{u}^{*}$ is the adjoint of a composition, we also have that $\dot{\zeta}_{u}^{*}(w)=\dot{\zeta}_{1,u}^{*}\circ\dot{\zeta}_{2,\zeta_{1}(u)}^{*}(w)$. Plugging in the values for the differential operators and adjoints of $\zeta_{1},\zeta_{2}$, this yields that $\dot{\zeta}_{u}(t)(x)=t(1,x)$ and $\dot{\zeta}_{u}^{*}(w)(a,x)=\frac{a}{\pi(x)}w(x)$.
C.5   Primitives for marginal quantities and semiparametric models
@@
The primitives in Table 1 can be adapted to a semiparametric model. This can be done by using that, for any semiparametric $\mathcal{M}^{\prime}\subset\mathcal{M}$, the restriction of a totally pathwise differentiable primitive $\theta:\mathcal{M}\times\mathcal{U}\rightarrow\mathcal{W}$ to ×$\mathcal{M}^{\prime}\times\mathcal{U}$ is itself totally pathwise differentiable. The adjoint of this new primitive at $(P,u)$ is the same as that of $\theta$, except its first entry is replaced by its orthogonal projection onto the tangent space ˙$\dot{\mathcal{M}}_{P}^{\prime}$. Thus, when $\psi:\mathcal{M}^{\prime}\rightarrow\mathbb{R}$ can be expressed as a composition of restrictions of primitives in Table 1, Algorithm 2 returns its semiparametric EIF relative to $\mathcal{M}^{\prime}$. Naturally, the returned form of the EIF will be easiest to work with when the projection onto ˙$\dot{\mathcal{M}}_{P}^{\prime}$ has a known closed form. While this is the case in some settings (e.g., Bickel et al., 1993, Chapter 3.2), it is certainly not the case for a general infinite-dimensional model. Identifying a means to circumvent this projection is an interesting area for future work.
Appendix D Step-by-step illustrations of Algorithm 2
Like Table 2 from the main text, Tables S1 and S2 provide step-by-step illustrations of Algorithm 2. The first studies the expected density parameter, and the second studies an expected conditional covariance parameter. The primitives used to express these parameters are detailed in the table captions. The value of $f_{0}$ in the bottom row of each table is the EIF at a distribution $P$ in a nonparametric model $\mathcal{M}$. In these examples, $Q\in\mathcal{M}$ is a fixed distribution used to define the ambient Hilbert spaces of some of the primitives used to represent $\psi$. We suppose $Q$ is mutually absolutely continuous with $P$ and let $\frac{p}{q}:=\frac{dP}{dQ}$ and $\frac{p_{X}}{q_{X}}:=\frac{dP_{X}}{dQ_{X}}$.
Table S1: Step-by-step evaluation of Algorithm 2’s computations of the EIF of the expected density $\psi(\tilde{P})=\int\tilde{p}(z)^{2}\lambda(dz)$, where $\tilde{p}=d\tilde{P}/d\lambda$ for a $\sigma$-finite measure $\lambda$. Under regularity conditions, the parameter $\psi$ is expressible as a composition of a root-density $\theta_{1}(\tilde{P},0)=\tilde{p}^{1/2}\in L^{2}(\lambda)$, pointwise square $\theta_{2}(\tilde{P},h_{1})=h_{1}^{2}\in L^{2}(\lambda)$, change of measure $\theta_{3}(\tilde{P},h_{2})=h_{2}\in L^{2}(Q)$, and mean operator $\theta_{4}(\tilde{P},h_{3})=E_{\tilde{P}}[h_{3}(Z)]\in\mathbb{R}$. Here, $\mathrm{pa}(1)=\emptyset$, $\mathrm{pa}(2)=\{1\}$, $\mathrm{pa}(3)=\{2\}$, and $\mathrm{pa}(4)=\{3\}$. The value of $f_{0}$ in the bottom row is the efficient influence function of $\psi$ at $P$.
@@ 0 $\bm{2[p-\psi(P)]}$
This lemma is a direct consequence of Lemma S5 and Appendix C.3.1 and so the proof is omitted.
Table S2: Step-by-step evaluation of Algorithm 2’s computations of the EIF of the expected conditional covariance $\psi(\tilde{P})=E_{\tilde{P}}[\mathrm{Cov}_{\tilde{P}}(A,Y\,|\,X)]$, where $Z=(X,A,Y)$. Let $\mathscr{C}:z\mapsto x$. Under regularity conditions, the parameter $\psi$ is expressible as a composition of conditional means $\theta_{1}(\tilde{P},0)(\cdot)=E_{\tilde{P}}[Y\,|\,X=\cdot\,]\in L^{2}(Q_{X})$ and $\theta_{2}(\tilde{P},0)(\cdot)=E_{\tilde{P}}[A\,|\,X=\cdot\,]\in L^{2}(Q_{X})$, liftings $\theta_{3}(\tilde{P},h_{1})=(z\mapsto h_{1}(x))\in L^{2}(Q)$ and $\theta_{4}(\tilde{P},h_{2})=(z\mapsto h_{2}(x))\in L^{2}(Q)$, a pointwise operation $\theta_{5}(\tilde{P},(h_{3},h_{4}))=(z\mapsto ay-h_{3}(z)h_{4}(z))\in L^{2}(Q)$, and a mean operator $\theta_{6}(\tilde{P},h_{5})=E_{\tilde{P}}[h_{5}(Z)]\in\mathbb{R}$. Here, $\mathrm{pa}(1)=\mathrm{pa}(2)=\emptyset$, $\mathrm{pa}(3)=\{1\}$, $\mathrm{pa}(4)=\{2\}$, $\mathrm{pa}(5)=\{3,4\}$, and $\mathrm{pa}(6)=\{5\}$. The value of $f_{0}$ in the bottom row is the efficient influence function of $\psi$ at $P$, where $\mu_{P}^{Y}(x):=E_{P}[Y\,|\,X=x]$ and $\mu_{P}^{A}(x):=E_{P}[A\,|\,X=x]$. Other expressions of this parameter, such as ones using the conditional covariance primitive studied in Appendix C.2.3, can also be used to compute the EIF.
@@ 0 $\bm{z\mapsto[a-\mu_{P}^{A}(x)][y-\mu_{P}^{Y}(x)]-\psi(P)}$
Appendix E Study of remainder in von Mises expansion of nonparametric $R^{2}$
@@ 7 $-\frac{1}{\widehat{h}_{2}}$ 00 00 00 -- $\frac{\widehat{h}_{7}}{\widehat{h}_{2}^{2}}$- 00
Forward and backward pass nuisances are estimated as described in Appendix F. In the forward routines, this entails setting $\widehat{h}_{1}:z\mapsto y$, $\widehat{h}_{2}=\mathrm{Var}_{P_{\mathscr{I}}}(Y)$ with $P_{\mathscr{I}}$ the empirical distribution of $Z_{\mathscr{I}}:=\{Z_{i}:i\in\mathscr{I}\}$, $\widehat{h}_{3}:z\mapsto y$, $\widehat{h}_{4}:x\mapsto\widehat{E}_{4}[Y\,|\,X=x]$ with the conditional mean estimate returned by a nonparametric regression procedure, $\widehat{h}_{5}:z\mapsto\widehat{h}_{4}(x)$, $\widehat{h}_{6}:z\mapsto[y-\widehat{h}_{5}(z)]^{2}$, $\widehat{h}_{7}=E_{P_{\mathscr{I}}}[\widehat{h}_{6}(Z)]$, and $\widehat{h}_{8}=1-\widehat{h}_{7}/\widehat{h}_{2}$. In the backward routines, the only nuisances that require estimation are $E_{P}[Y\,|\,X=\,\cdot\,]$ and $E_{P}[Y]$, which arise when obtaining the estimates $\widehat{\vartheta}_{5}$ and $\widehat{\vartheta}_{2}$, respectively. We denote these estimates by $\widehat{E}_{5}[Y\,|\,X=\,\cdot\,]$ and $\widehat{E}_{2}[Y]$, respectively.
@@
Appendix F Nuisance estimation routines for primitives in Table 1
F.1   Overview
@@
F.2   Maps depend nontrivially on both of their arguments
F.2.1   Conditional mean